According to a 2009 Reader's Digest article, people throw away approximately 10% of what they buy at the grocery store. Assume this is the true proportion and you plan to randomly survey 153 grocery shoppers to investigate their behavior. What is the probability that the sample proportion exceeds 0.13?

Note: You should carefully round any z-values you calculate to 4 decimal places to match wamap's approach and calculations.

Answer =  (Enter your answer as a number accurate to 4 decimal places.)
Answer #1

Solution

Given that,

p = 0.10

1 - p = 1 - 0.10 = 0.90

n = 153

\mu\hat p = p = 0.10

\sigma\hat p =  \sqrt[p( 1 - p ) / n] = \sqrt [(0.10 * 0.90) / 153 ] = 0.0243

P( \hat p > 0.13) = 1 - P( \hat p < 0.13)

= 1 - P(( \hat p - \mu \hat p ) / \sigma \hat p < (0.13 - 0.10) / 0.0243)

= 1 - P(z < 1.2346)

Using z table

= 1 - 0.8915

= 0.1085